OpenStudy (lena772):

determine the point(s) at which the graph of the function has a horizontal tangent line f(x)=x^2/(x^2+1)

OpenStudy (anonymous):

( 0,0)

OpenStudy (lena772):

how did you solve that?

ganeshie8 (ganeshie8):

slope of horizontal line = 0

ganeshie8 (ganeshie8):

you can get slope by finding the derivative

ganeshie8 (ganeshie8):

so find the derivative of f(x) and see when it equals 0

OpenStudy (ybarrap):

When you find a derivative of a function (i.e. it's slope at each point), you are finding the tangents (i.e. slopes) at every point. Since the question is asking you to find "horizontal" tangents that equal zero, it's the same as asking you where the slope of this function is zero. Zero slopes for any function occur where minimums or maximums occur, which makes sense, because at these points, the slopes are going from negative to positive, so somewhere in-between there must be a zero slope. Do you know how to find derivatives?

OpenStudy (lena772):

Yes

OpenStudy (ybarrap):

Then let's do it, let's find the derivative of this function.

OpenStudy (lena772):

d/dx x^n=nx^n-1

OpenStudy (ybarrap):

$$ f(x)=\cfrac{x^2}{x^2+1}\\ f'(x)=? $$ Do you know how to use the quotient rule (bottom, times derivative of the top, minus the top, times the derivative of the bottom, divided by the bottom, squared)?

OpenStudy (ybarrap):

That formula is correct for a monomial, we need also to apply the quotient rule, using also this monomial derivative.

OpenStudy (ybarrap):

Let's do this in stops. What is the denominator multiplied times the derivative of the numerator?

OpenStudy (ybarrap):

*steps not stops lol

OpenStudy (ybarrap):

$$ f(x)=\cfrac{x^2}{x^2+1}\\ $$ Step 1 (bottom times derivative of the top): $$ (x^2+1)\times \cfrac{d}{dx}x^2=? $$

OpenStudy (lena772):

f(x)=x^2/x^2+1 f(x)= x^2/1 + x^2/x^2 f(x) = x^2+1

OpenStudy (ybarrap):

$$ \cfrac{x^2}{1}+\cfrac{x^2}{x^2}\ne\cfrac{x^2}{x^2+1}\\ $$ This is not equal because, $$ \cfrac{x^2}{1}+\cfrac{x^2}{x^2}=\cfrac{x^2}{1}\times\cfrac{x^2}{x^2}+\cfrac{x^2}{x^2}\\ =\cfrac{x^4}{1}+\cfrac{x^2}{x^2}=\cfrac{x^4+x^2}{x^2}\ne\cfrac{x^2}{x^2+1} $$

OpenStudy (ybarrap):

*correction to second part above: $$ \cfrac{x^2}{1}+\cfrac{x^2}{x^2}=\cfrac{x^2}{1}\times\cfrac{x^2}{x^2}+\cfrac{x^2}{x^2}\\ =\cfrac{x^4}{\color{red}{x^2}}+\cfrac{x^2}{x^2}=\cfrac{x^4+x^2}{x^2}\ne\cfrac{x^2}{x^2+1} $$

OpenStudy (lena772):

I'm lost now.

OpenStudy (ybarrap):

$$ \cfrac{1}{1+2}=\cfrac{1}{3}\ne\cfrac{1}{1}+\cfrac{1}{2}=\cfrac{3}{2}\\ \cfrac{1}{3}\ne\cfrac{3}{2} $$ This has nothing to do with your problem, but this is what you were starting off with and I want to make sure you understand, but because you need to know this.

OpenStudy (ybarrap):

Let's now do s step 1: $$ (x^2+1)\times \cfrac{d}{dx}x^2=? $$

OpenStudy (ybarrap):

$$ \cfrac{d}{dx}x^2=2x $$ So, $$ (x^2+1)\times \cfrac{d}{dx}x^2=(x^2+1)\times 2x=2x^3+2x $$ Make sense so far?

OpenStudy (lena772):

yes

OpenStudy (ybarrap):

That was bottom times derivative of the top. Step 2 (top times derivative of the bottom): $$ x^2\cfrac{d}{dx}(x^2+1)=? $$

OpenStudy (ybarrap):

$$ \cfrac{d}{dx}(x^2+1)=2x $$ So, $$ x^2\cfrac{d}{dx}(x^2+1)=x^2\times 2x=2x^3 $$ Are you okay with this?

OpenStudy (lena772):

Where does the 2x come from

OpenStudy (ybarrap):

We are doing the quotient rule: Step 1: bottom times derivative of the top Step 2: top times derivative of the bottom Step 3: bottom squared The we combine the results of each: $$ \cfrac{\text{Step 1 - Step 2}}{{\text{Step 3}^2}} $$

OpenStudy (ybarrap):

The 2x comes from this: $$ \cfrac{d}{dx}(x^2+1)=2x $$ This is using your original formula for the derivative of a monomial: d/dx x^n=nx^n-1. Where here, n=2.

OpenStudy (ybarrap):

$$ \cfrac{d}{dx}(x^2+1)=\cfrac{d}{dx}x^2+\cfrac{d}{dx}1=2\times x^{2-1}+0=2x $$

OpenStudy (ybarrap):

questions?

OpenStudy (ybarrap):

$$ f(x)=\cfrac{x^2}{x^2+1}\\ f'(x)=0\implies\cfrac{(x^2+1)2x-x^2\times2x}{\left (x^2+1\right )^2}=\cfrac{2x}{x^4+2x^2+1}=0\implies x=0 $$ So, horizontal tangents occur at \(x=0\). Answer: At \(x=0\): \(f(0)=\cfrac{0^2}{0^2+1}=0\). So, \(f(x)=0\) at the horizontal tangents.